In year 2004 the person weighted 160 pounds, year 2005 the person weighed 168 pounds, and year 2006 the person weighed 175 pounds, create an equation or inequality of the year and the weight. Draw a straight line graph to depict the variable in the situation of the weight increases.

Answer by stanbon(75887)

In year 2004 the person weighted 160 pounds, year 2005 the person weighed 168 pounds, and year 2006 the person weighed 175 pounds, create an equation or inequality of the year and the weight. Draw a straight line graph to depict the variable in the situation of the weight increases.
Let 4 represent 2004 and 5 represen 2005
You have two points: (4,160) and (5,175)
Slope = [175-160]/[5-1]=15
Since y=mx+b, 160=15(4)+b
b=160-60=100
EQUATION:
weight = 15(year)+100
Use the two points or the equation to graph the relation.
Cheers,
stan H.